OpenStudy (anonymous):

Which is the first negative term in this sequence? 21, 19, 17 . . . 7th term 8th term 9th term 12th term

OpenStudy (solomonzelman):

what do you do each time to get to the next term ?

OpenStudy (anonymous):

go down 2

OpenStudy (solomonzelman):

yes, you subtract 2.

OpenStudy (solomonzelman):

Now, every term (whether positive or negative) is going to be ODD. Agree with this?

OpenStudy (anonymous):

12th term

OpenStudy (anonymous):

did you think of 12 when you first came cause you started answering pretty fast

OpenStudy (anonymous):

or did you just knew you could answer this answer lol

OpenStudy (solomonzelman):

12th term? \(\large\color{black}{ \displaystyle -1=21+(-2)(n-1) }\) \(\large\color{black}{ \displaystyle -1=21-2n+2 }\) \(\large\color{black}{ \displaystyle -3=21-2n }\) \(\large\color{black}{ \displaystyle -24=-2n }\) \(\large\color{black}{ \displaystyle n=12 }\)

OpenStudy (solomonzelman):

yes, 12th

OpenStudy (solomonzelman):

We used \(\large\color{black}{ \displaystyle a_n=a_1+d(n-1) }\)

OpenStudy (solomonzelman):

and here we answered the question "whichTH term in the sequence is -1 ? "

OpenStudy (solomonzelman):

Why -1? because it is the first negative term.

OpenStudy (solomonzelman):

again, correct the first negative term is the 12th...

OpenStudy (anonymous):

thx